OpenStudy (anonymous):

What is the solution set for |–x + 2| = 5? x = –7 and x = –3 x = –7 and x = 3 x = 7 and x = –3 x = 7 and x = 3

terenzreignz (terenzreignz):

The thing about absolute values is that it alters the way you solve your equations. The key thing here is given |a| = b It could mean two things, either a = b OR a = -b

OpenStudy (anonymous):

ohkay,...

terenzreignz (terenzreignz):

So in your example you have |-x + 2| = 5 So, as stated above, you have consider both -x + 2 = 5 OR -x + 2 = -5 Try this now...

OpenStudy (anonymous):

uhm,... I dont get it

terenzreignz (terenzreignz):

You get why if |a| = b it could mean either a = b or a = -b right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x could be positive number or negative number

terenzreignz (terenzreignz):

Well, think of -x + 2 as your a and 5 as your b

OpenStudy (anonymous):

nicely explained @ terenzreignz

terenzreignz (terenzreignz):

Thanks :) @miah Stuck?

terenzreignz (terenzreignz):

Well, as shown, given |a| = b the way to do it, provided b is positive, is to do this: a = b and a = -b Catch me so far?

OpenStudy (anonymous):

@terenzreignz show an example so maybe she would understand

terenzreignz (terenzreignz):

Excellent suggestion :) Hey, @miah Suppose we have 2|x - 8|=6 And we're to find the possible values for x.

terenzreignz (terenzreignz):

First, we divide both sides by 2, to get rid of that annoying coefficient. We get |x - 8| = 3 Catch me so far?

terenzreignz (terenzreignz):

So, this means either x - 8 = 3 OR x - 8 = -3 We solve both equations, we get x = 11 and x = 5 And these are the answers.